OpenStudy (abbycross167):

Can someone please show me the steps to answer this question? If x = 2 and y = -3, what is the value of 2x^2 - 3xy – 2y^2

OpenStudy (vhtran):

Just plug in your X and Y values

OpenStudy (abbycross167):

so 22^2 - 32-3– 2-3^2

OpenStudy (vhtran):

2(2\[2(2^2)-3(2)(-3)-2(-3^2)\]

OpenStudy (vhtran):

That should be right

OpenStudy (vhtran):

After that just simplify

OpenStudy (abbycross167):

So 2(2^2) - 3(2)(-3) – 2(2^2) = 2(4) - 3(2)(-3) – 2(4)

OpenStudy (vhtran):

yup

OpenStudy (vhtran):

then multiply

OpenStudy (vhtran):

2X4-3X-6-8

OpenStudy (abbycross167):

So 2(2^2) - 3(2)(-3) – 2(2^2) = 2(4) - 3(2)(-3) – 2(4) = 2*4=8 3*2=6 6*-3=-18 2*4=8 =

OpenStudy (abbycross167):

8-18-8?

OpenStudy (vhtran):

8--18-8, Don'y forget the negative on the 18 :P

OpenStudy (vhtran):

What happens when you have 2 negatives?

OpenStudy (abbycross167):

positive right?

OpenStudy (vhtran):

yes :D so 8+18-8

OpenStudy (abbycross167):

= 18

OpenStudy (mathstudent55):

You need to be careful when you square a negative number. \(-3^2 \ne (-3)^2\) \(-3^2 = -9\) \((-3)^2 = 9\) Also, make sure you substitute the correct numbers in for the variables. \(\color{red}{x = 2}\), and \(\color{green}{y = -3}\). \( 2\color{red}{x}^2 - 3\color{red}{x}\color{green}{y} – 2\color{green}{y}^2 \) \(= 2(\color{red}{2})^2 - 3(\color{red}{2})(\color{green}{-3}) - 2(\color{green}{-3})^2\) \(=2(4) - (-18) - 2(9)\) \(= 8 + 18 - 18\) \(= 8\)

OpenStudy (mathstudent55):

The correct answer is 8, not 18.

OpenStudy (mathstudent55):

Do you understand the steps?

OpenStudy (abbycross167):

Yes sir/ma'am, thank you very very much!!

OpenStudy (mathstudent55):

You're welcome.